# revtex4-2Repair the float

### v1

@@ II.2 Many Parameter Learning
Figure 9: Difference in 3 quantum dot in charge stability diagrams for input and expected measurement outcomes from Hubbard model parameters for the prediction of disorder in all parameters. In particular $n_{error}=||\langle\vec{n}_{input}\rangle-\langle\vec{n}_{predicted}\rangle||$ is plotted in both plots where $\langle\vec{n}_{input}\rangle$ is the occupation expectation vector for all sites of the input data and $\langle\vec{n}_{predicted}\rangle$ is the expected occupation expectation vector given the prediction of the Hubbard model parameters. The columns of both plots correspond to the charge stability diagrams in Fig. 7 with the same column. The error-free model parameters were set to $U=4$, $t=1$ and $V_{i,i+1}=0.2$. (a) $n_{error}$ between input Hubbard parameters and predicted Hubbard parameters where $\mu_{1}$ and $\mu_{3}$ are independently varied. (b) $n_{error}$ between input Hubbard parameters and predicted Hubbard parameters where the chemical potential at each site is $\vec{\mu}=[\mu_{1},\cdots,\mu_{n}]$ with our plot having axis $\mu_{1}=-\mu_{3}$ (rescaled by $\langle U\rangle$ and shifted by $\mu$) vs. $\mu_{k\neq 1,3}=\mu$ (rescaled by $W=\frac{1}{L}\left(\sum_{i}\langle U_{i}\rangle+\langle V_{i,i+1}\rangle\right)$ where $\langle U\rangle=4$ and $\langle V_{i,i+1}\rangle=0.2$ are the their disorder free values respectively).
Figure 10: 4 quantum dot charge stability diagrams for input and expected measurement outcomes from Hubbard model parameters for the prediction of disorder in all parameters. The root mean squared error in the Hubbard model parameters was $R_{MS}({\delta\epsilon_{i}})=0.0280$, $R_{MS}(\delta V_{i,j})=0.0205$, $R_{MS}(\delta t_{i,j})=0.0193$, $R_{MS}(\delta U_{i})=0.0512$ with an $R^{2}=0.9961$. In both plots the first row is the input charge stability diagrams, namely the most probable state for a chemical potential configuration, and the second row is the expected charge stability diagram given the prediction of the Hubbard model parameters. (a) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where $\mu_{1}$ and $\mu_{3}$ are independently varied. (b) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where the chemical potential at each site is $\vec{\mu}=[\mu_{1},\cdots,\mu_{n}]$ with our plot having axis $\mu_{1}=-\mu_{3}$ (rescaled by $\langle U\rangle$ and shifted by $\mu$) vs. $\mu_{k\neq 1,3}=\mu$ (rescaled by $W=\frac{1}{L}\left(\sum_{i}\langle U_{i}\rangle+\langle V_{i,i+1}\rangle\right)$ where $\langle U\rangle=4$ and $\langle V_{i,i+1}\rangle=0.2$ are the their disorder free values respectively ), this is similar to the stability diagrams fed into the neural network except the neural network only receives $\mu_{i}=-\mu_{j}$ between nearest neighbors $i=j+1$.
Figure 12: Difference in 4 quantum dot in charge stability diagrams for input and expected measurement outcomes from Hubbard model parameters for the prediction of disorder in all parameters. In particular $n_{error}=||\langle\vec{n}_{input}\rangle-\langle\vec{n}_{predicted}\rangle||$ is plotted in both plots where $\langle\vec{n}_{input}\rangle$ is the occupation expectation vector for all sites of the input data and $\langle\vec{n}_{predicted}\rangle$ is the expected occupation expectation vector given the prediction of the Hubbard model parameters. The columns of both plots correspond to the charge stability diagrams in Fig. 10 with the same column. The error-free model parameters were set to $U=4$, $t=1$ and $V_{i,i+1}=0.2$. (a) $n_{error}$ between input Hubbard parameters and predicted Hubbard parameters where $\mu_{1}$ and $\mu_{3}$ are independently varied. (b) $n_{error}$ between input Hubbard parameters and predicted Hubbard parameters where the chemical potential at each site is $\vec{\mu}=[\mu_{1},\cdots,\mu_{n}]$ with our plot having axis $\mu_{1}=-\mu_{3}$ (rescaled by $\langle U\rangle=4$ and shifted by $\mu$) vs. $\mu_{k\neq 1,3}=\mu$ (rescaled by $W=\frac{1}{L}\left(\sum_{i}\langle U_{i}\rangle+\langle V_{i,i+1}\rangle\right)$ where $\langle U\rangle=4$ and $\langle V_{i,i+1}\rangle=0.2$ are the their disorder free values respectively)
@@ II.3 No Trust Verification
Here, we simulate performing measurements in these different configurations (including new measurements beyond occupation numbers) to generate a new $X^{\prime}$. Since one could then carry out those measurements experimentally for the $K^{\prime}$ measurement configurations and compare their outcomes to that of $X^{\prime}$, a match would serve as strong evidence of the validity of the disorder (or, more specifically, Hubbard model parameter) predictions. One could generate an arbitrarily large set of new measurement configurations, $K^{\prime}$, to test the predictions, and all reasonable $K$ matrices should still match in both simulated and experimental results if the $Y$ prediction is accurate. Since this new $K^{\prime}$ never enters the machine learning scheme, the neural network would be incapable of generating Hubbard parameters that "hallucinate" a match with these measurements. Thus, without any trust of the machine learning scheme, one could perform as many measurements as necessary to ensure that the Hubbard model parameters predicted by the algorithm are correct.
@@ II.4 Costs
Let us consider how this method would be applied to a large number of quantum dots. It is important to note that our method cannot be scaled up indefinitely, as eventually, the exponential scaling of the Hubbard model simulations will render it unfeasible. (This has nothing whatsoever to do with any limitation in our machine learning method or algorithm, it is simply that generating Hubbard model exact diagonalization solutions is an exponentially hard problem, which becomes unfeasible for more than 10-20 quantum dots because of the exponential growth of the Hilbert space.) However, it does permit simultaneous tuning of many more dots in aggregate at once than manual tuning allows. It is crucial to emphasize that the machine learning aspect is not the limiting factor; in fact, we find that the neural network and the amount of training data required do not seem to scale significantly with the addition of more quantum dots. The primary limitation is the exponential scaling of the Hubbard model used in generating the training data, which restricts our method’s scalability to an arbitrary number of quantum dots. The fact that the machine learning numerics itself do not scale with the number of dots indicates that perhaps our technique may be useful in quantum computing platforms with hundreds of quantum dot qubits since automating the control may involve only 2-5 dots at a time, as discussed below.
@@
The notion of scaling up to tune an arbitrary number of quantum dots simultaneously is, to begin with, excessive and unnecessary. The necessity for tuning quantum dots simultaneously is not dictated by the total number of dots within the system but by the locality of the system. For a large system with, say, 100s or more quantum dots, the envisioned tuning process involves adjusting a quantum dot within a locality that includes a certain number of nearest and next-nearest neighbors dependent on the interaction range. In this localized window within the bulk, one would determine the parameters of the central quantum dot(s) and then shift the window through the system, tuning a limited number of quantum dots at a time while considering the effects of a broader locality on the charge stability diagram. It is worth mentioning that as the window shifts, a new neural network would not need to be trained; one could simply reuse the charge stability diagrams measured from the initial locality for subsequent tunings. Thus, our technique may provide the pathway for the automated control of Hubbard model based quantum dot platforms of the future containing arbitrarily large number of qubits.
The ability of our method to tune more quantum dots simultaneously becomes particularly crucial when considering 2D geometries. In a 2D grid of quantum dots, current methods that tune two or three quantum dots at the same time are likely to face challenges, as each dot would have four nearest neighbors, even if we only consider nearest neighbors as relevant. In such a case even these nearest neighbors would significantly influence the charge stability diagram between any two dots. Our method, in principle, could be adapted to tackle such scenarios, enabling the tuning of much more complex geometries through the use of a specified locality window. Moreover, since current systems utilize quantum dot chains, it appears that tuning up to 5 quantum dots at a time, as we currently achieve, would be sufficient for a system with any number of dots since the interactions are likely to be local and limited to a few (<5) dots only.
@@ III Conclusion
In conclusion, we provide a method of solving disorder landscapes though a series of aggregate measurements that could be generalized to many physical models. In this work we also provide a novel "no trust" verification method that virtually eliminates the risk of "hallucination" predictions by the neural network. This provides a method of conclusively verifying the predictions of the neural network such that even without any trust of the neural network one can be confident about the results.
@@
This work successfully demonstrates a method, which by employing machine learning techniques, specifically through the use of a CNN, is able to accurately determine disordered deviations in the parameters of the extended Hubbard model as appied to quantum dot qubits. By analyzing a series of charge stability diagrams from nearest neighbor pairs of quantum dots, our CNN is adept at identifying spatially dependent disorder across each parameter of the model. This approach shows remarkable proficiency in predicting with high accuracy the spatially dependent variations in coupling constants, both Coulomb intra and inter-site terms, as well as site-specific gate voltage errors. These results are of particular experimental interest due to the natural description of coupled semiconductor quantum dots through the Hubbard model and the readily experimental availability of charge stability diagram measurements. Although we discuss our results in terms of disorder, our technique in fact is a method to obtain the extended Hubbard model Hamiltonian parameters themselves using just the stability diagrams as the input. Even if all the parameters underlying the quantum dot qubit platform are unknown, our technique, in principle, can provide an accurate estimate for all the parameters of the underlying Hamiltonian as long as sufficient amount of input charge stability diagrams are available. Thus, the technique developed in this work is not only applicable to semiconductor quantum dot qubit platforms, but also to the strongly correlated Hubbard model emulations.
@@
Extending beyond previous studies, our investigation includes the prediction of disorder not only in individual parameters per site but also across all parameters simultaneously. This provides a comprehensive understanding of disorder within the system, marking a notable enhancement over earlier methods that focused primarily on single parameter deviations, such as gate voltages [10]. Additionally, our research broadens the scope of conventional quantum dot tuning practices by considering the interactions among 3, 4, and 5 quantum dots simultaneously, thereby integrating crosstalk effects that are typically disregarded in the standard methodology of tuning quantum dot pairs.
This study is the first to solve site-specific disordered parameters within the extended Hubbard model for configurations exceeding two quantum dots, covering aspects such as coupling constants, on-site potentials, and inter-site repulsions, and extending to site-specific disordered gate voltages for more than three quantum dots. The introduction of our CNN not only paves the way for potentially fully automated tuning of quantum dot-based systems, eliminating the laborious manual adjustment process when parameters drift, but also ensures the reliability of the tuning process through a "no trust" verification method.
@@
We mention that in quantum dot qubits (and in fact, in all quantum computing platforms), the background disorder (which is the topic of the current work) drifts in the sample changing the disorder landscape slowly over time, creating a charge noise. Typically, these drifts are slow and the noise is quasistatic, which is what our model for disorder assumes. But such a drift of course causes the practical problem that an already perfectly tuned experimental quantum dot system drifts out of its tuned parameter regime over time, and must be retuned. This is the key problem. Our machine learning technique developed in this work, when embedded into the system, would enable the retuning in an automated way without manual human intervention since all that is needed is a measurement of the charge stability diagram allowing the system to be retuned, which can be incorporated into the measurement and tuning electronics. The only necessary ingredient is the training of the algorithm using a sufficient amount of input stability diagrams obtained from measurements on the quantum dot circuits, as established in the current work using our simulations with different background disorder (which simulates the effective drift and noise).
@@
Further, though the use of higher resolution charge stability diagram input, our investigation seems to show that for smaller coupling constants the accuracy in predictions of the disorder can be increased as as much as necessary to fit the needs of experimentalists. We also show that a smaller $t_{ij}$ tends to increase the fidelity of the results in the other parameters. Additionally, our results show that the method continues to work as the number of quantum dots increases with little to no decrease in fidelity. While we do not envision using our method for a whole computational device (due to the scaling of computation in generating the training data), our technique could be scaled up to match the locality of a device to tune many quantum dots at once.
@@
Looking to the future, while our method is currently able to apply directly to experiment and significantly improve the present tuning process of quantum dot devices, it should be possible to extend it to start from a microscopic Hamiltonian to determine directly the semiconductor quantum dot parameters. In fact, there is no reason to think that the technique itself is limited only to the Hubbard model, and in principle, it can be used to study other models of quantum dot qubits effectively. Our current use of the extended Hubbard model is based only on the fact that such a model seems to describe the semiconductor quantum dot based qubit platforms reasonably well.
@@ IV Acknowledgement
Figure 15: 4 quantum dot charge stability diagrams for input and expected from Hubbard model parameters for the prediction of only $\delta V_{ij}$. The root mean squared error in the Hubbard model $\delta V_{ij}$ parameter was $R_{MS}({\delta V_{ij}})=$ with an $R^{2}=0.99995$. In both plots the first row is the input charge stability diagrams, namely the most probable state for a chemical potential configuration, and the second row is the expected charge stability diagram given the prediction of the Hubbard model parameters. The error-free model parameters were set to $U=4$, $t=1$ and $V_{i,i+1}=0.2$. The columns of these two subplots correspond to the same representative samples. (a) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where $\mu_{1}$ and $\mu_{3}$ are independently varied. (b) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where the chemical potential at each site is $\vec{\mu}=[\mu_{1},\cdots,\mu_{n}]$ with our plot having axis $\mu$ (rescaled by $\langle U\rangle$ and shifted by $\mu$) vs. $\mu_{k\neq 1,3}=\mu$ (rescaled by $W=\frac{1}{L}\left(\sum_{i}\langle U_{i}\rangle+\langle V_{i,i+1}\rangle\right)$ where $\langle U\rangle=4$ and $\langle V_{i,i+1}\rangle=0.2$ are the their disorder free values respectively), this is similar to the stability diagrams fed into the neural network except the neural network only receives $\mu_{i}=-\mu_{j}$ between nearest neighbors $i=j+1$.
Figure 16: 4 quantum dot charge stability diagrams for input and expected from Hubbard model parameters for the prediction of only $\delta t_{ij}$. The root mean squared error in the Hubbard model parameter was $R_{MS}({\delta t_{ij}})=0.0243$ with an $R^{2}=0.99997$. In both plots the first row is the input charge stability diagrams, namely the most probable state for a chemical potential configuration, and the second row is the expected charge stability diagram given the prediction of the Hubbard model parameters. The error-free model parameters were set to $U=4$, $t=1$ and $V_{i,i+1}=0.2$. The columns of these two subplots correspond to the same representative samples. (a) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where $\mu_{1}$ and $\mu_{3}$ are independently varied. (b) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where the chemical potential at each site is $\vec{\mu}=[\mu_{1},\cdots,\mu_{n}]$ with our plot having axis $\mu_{1}=-\mu_{3}$ (rescaled by $\langle U\rangle$ and shifted by $\mu$) vs. $\mu_{k\neq 1,3}=\mu$ (rescaled by $W=\frac{1}{L}\left(\sum_{i}\langle U_{i}\rangle+\langle V_{i,i+1}\rangle\right)$ where $\langle U\rangle=4$ and $\langle V_{i,i+1}\rangle=0.2$ are the their disorder free values respectively ), this is similar to the stability diagrams fed into the neural network except the neural network only receives $\mu_{i}=-\mu_{j}$ between nearest neighbors $i=j+1$.
Figure 17: 4 quantum dot charge stability diagrams for input and expected from Hubbard model parameters for the prediction of only $\delta U_{i}$. The root mean squared error in the Hubbard model $\delta U_{i}$ parameter was $R_{MS}({\delta U_{i}})=0.1005$ with an $R^{2}=0.9987$. In both plots the first row is the input charge stability diagrams, namely the most probable state for a chemical potential configuration, and the second row is the expected charge stability diagram given the prediction of the Hubbard model parameters. The error-free model parameters were set to $U=4$, $t=1$ and $V_{i,i+1}=0.2$. The columns of these two subplots correspond to the same representative samples. (a) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where $\mu_{1}$ and $\mu_{3}$ are independently varied. (b) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where the chemical potential at each site is $\vec{\mu}=[\mu_{1},\cdots,\mu_{n}]$ with our plot having axis $\mu_{1}=-\mu_{3}$ (rescaled by $\langle U\rangle$ and shifted by $\mu$) vs. $\mu_{k\neq 1,3}=\mu$ (rescaled by $W=\frac{1}{L}\left(\sum_{i}\langle U_{i}\rangle+\langle V_{i,i+1}\rangle\right)$ where $\langle U\rangle=4$ and $\langle V_{i,i+1}\rangle=0.2$ are the their disorder free values respectively), this is similar to the stability diagrams fed into the neural network except the neural network only receives $\mu_{i}=-\mu_{j}$ between nearest neighbors $i=j+1$.
Figure 18: 3 quantum dot charge stability diagrams for input and expected measurement outcomes from Hubbard model parameters for the prediction of disorder in all parameters with higher resolution, standard scaled, and reduced range $\tilde{t}\in[0.01,0.25]$. The root mean squared error in the Hubbard model parameters was $R_{MS}({\delta\epsilon_{i}})=0.0110$, $R_{MS}({\delta V_{i,j}})=0.0113$, $R_{MS}({\delta t_{i,j}})=0.0045$, $R_{MS}({\delta U_{i}})=0.0347$ with an $R^{2}=0.9972$. In both plots the first row is the input charge stability diagrams, namely the most probable state for a chemical potential configuration, and the second row is the expected charge stability diagram given the prediction of the Hubbard model parameters. (a) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where $\mu_{1}$ and $\mu_{3}$ are independently varied. (b) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where the chemical potential at each site is $\vec{\mu}=[\mu_{1},\cdots,\mu_{n}]$ with our plot having axis $\mu_{1}=-\mu_{3}$ (rescaled by $\langle U\rangle$ and shifted by $\mu$) vs. $\mu_{k\neq 1,3}=\mu$ (rescaled by $W=\frac{1}{L}\left(\sum_{i}\langle U_{i}\rangle+\langle V_{i,i+1}\rangle\right)$ where $\langle U\rangle=4$ and $\langle V_{i,i+1}\rangle=0.2$ are the their disorder free values respectively), this is the data input into the machine learning model.

### v2

@@ II.3 No Trust Verification Scheme
Here, we propose to simulate performing measurements in these different configurations (including new measurements beyond occupation numbers) to generate a new $X^{\prime}$. Since one could then carry out those measurements experimentally for the $K^{\prime}$ measurement configurations and compare their outcomes to that of $X^{\prime}$, a match would serve as strong evidence of the validity of the disorder (or, more specifically, Hubbard model parameter) predictions. One could generate an arbitrarily large set of new measurement configurations, $K^{\prime}$, to test the predictions, and all reasonable $K$ matrices should still match in both simulated and experimental results if the $Y$ prediction is accurate. Since this new $K^{\prime}$ never enters the machine learning scheme, the neural network would be incapable of generating Hubbard parameters that "hallucinate" a match with these measurements. Since the probability of a "hallucination" result passing these additional verification checks is low, it would likely not be necessary to perform many additional measurements. Thus, without any trust of the machine learning scheme, one could perform as many measurements as necessary to ensure that the Hubbard model parameters predicted by the algorithm are correct. These verifications also may allow additional experimental tuning of the NN in the cases of failure, for instance minor modifications to the model Hamiltonian, a higher resolution in the charge stability diagram or merely more training data could be required to achieve accuracy for some experimental applications, the verification scheme can work to do this without being mislead by potential "hallucinations".
@@ II.4 Costs
Let us consider how this method would be applied to a large number of quantum dots. It is important to note that our method cannot be scaled up indefinitely, as eventually, the exponential scaling of the Hubbard model simulations will render it unfeasible. (This has nothing whatsoever to do with any limitation in our machine learning method or algorithm, it is simply that generating Hubbard model exact diagonalization solutions is an exponentially hard problem, which becomes unfeasible for more than 15-20 quantum dots because of the exponential growth of the Hilbert space.) However, it does permit simultaneous tuning of many more dots in aggregate at once than manual tuning allows. It is crucial to emphasize that the machine learning aspect is not the limiting factor; in fact, we find that the neural network and the amount of training data required do not seem to scale significantly with the addition of more quantum dots. The primary limitation is the exponential scaling of the Hubbard model used in generating the training data, which restricts our method’s scalability to an arbitrary number of quantum dots.
@@
The notion of scaling up to tune an arbitrary number of quantum dots simultaneously is, to begin with, excessive and unnecessary. The necessity for tuning quantum dots simultaneously is not dictated by the total number of dots within the system but by the locality of the system. For a large system with, say, 100s or more quantum dots, the envisioned tuning process involves adjusting a quantum dot within a locality that includes a certain number of nearest and next-nearest neighbors dependent on the interaction range. The effect another dot has on the central dots model parameters decreases the further away it is. In this localized window within the bulk, one would determine the parameters of the central quantum dot(s) and then shift the window through the system, tuning a limited number of quantum dots at a time while considering the effects of a broader locality on the charge stability diagram. It is worth mentioning that as the window shifts, a new neural network would not need to be trained; one could simply reuse the charge stability diagrams measured from the initial locality for subsequent tunings. Thus, our technique may provide the pathway for the automated control of Hubbard model based quantum dot platforms of the future containing arbitrarily large number of qubits.
The ability of our method to tune more quantum dots simultaneously becomes particularly crucial when considering 2D geometries. In a 2D grid of quantum dots, current methods that tune two or three quantum dots at the same time are likely to face challenges, as each dot would have four nearest neighbors, even if we only consider nearest neighbors as relevant. In such a case even these nearest neighbors would significantly influence the charge stability diagram between any two dots. Our method, in principle, could be adapted to tackle such scenarios, enabling the tuning of much more complex geometries through the use of a specified locality window. Moreover, in terms of linear quantum dot chains, we suspect 5 dots may be sufficient for most near-term devices since using current manual methods chains of length 12 are able to be tuned by 2-3 dots at a time [19]. In terms of scaling to 100s of dots with complex geometries, while the locality required is not known and thus our method is not guaranteed to work for arbitrarily large arrays, it is a potentially viable path for which no alternative exists.
@@ III Conclusion
In conclusion, we provide a method of solving disorder landscapes though a series of aggregate measurements that could be generalized to many physical models. In this work we also propose a novel "no trust" verification method that virtually eliminates the risk of "hallucination" predictions by the neural network. This provides a method of conclusively verifying the predictions of the neural network such that even without any trust of the neural network one can be confident about the results.
@@
Our technique explicitly uses deep learning on the whole system together and is not focused on tuning small parts (e.g. a few dots in a many quantum dot system) of the system individually using automated control, making our method particularly useful and powerful. This work successfully demonstrates a method, which by employing machine learning techniques, specifically through the use of a CNN, is able to accurately determine disordered deviations in the parameters of the extended Hubbard model as applied to quantum dot qubits. By analyzing a series of charge stability diagrams from nearest neighbor pairs of quantum dots, our CNN is adept at identifying spatially dependent disorder across each parameter of the model. This approach shows remarkable proficiency in predicting with high accuracy the spatially dependent variations in coupling constants, both Coulomb intra and inter-site terms, as well as site-specific gate voltage errors. These results are of experimental interest due to the natural description of coupled semiconductor quantum dots through the Hubbard model and the readily experimental availability of charge stability diagram measurements. Although we discuss our results in terms of disorder, our technique in fact is a method to obtain the extended Hubbard model Hamiltonian parameters themselves using just the stability diagrams as the input. Even if all the parameters underlying the quantum dot qubit platform are unknown, our technique, in principle, can provide an accurate estimate for all the parameters of the underlying Hamiltonian as long as sufficient amount of input charge stability diagrams are available. This is without the need for any additional information on the system, beyond some rough (potentially order of magnitude) estimate on what values the Hubbard parameters can take for training data generation. We emphasize, it is not necessary to get an initial site specific estimate of any Hubbard model parameter, a vague all sites have $U\in[0,8]$ for instance is sufficient to generate training data and predict any particular sites $\tilde{U}_{i}$. Thus, the technique developed in this work is not only applicable to semiconductor quantum dot qubit platforms, but also to the strongly correlated Hubbard model emulations.
@@
Extending beyond previous studies, our investigation includes the prediction of disorder not only in individual parameters per site but also across all parameters simultaneously. This provides a comprehensive prediction of disorder within the system, marking a notable enhancement over earlier methods that focused primarily on single parameter deviations, such as gate voltages [14]. Additionally, our research broadens the scope of conventional quantum dot tuning practices by considering the interactions among 3, 4, and 5 quantum dots simultaneously, thereby integrating crosstalk effects that are typically disregarded in the standard methodology of tuning quantum dot pairs.
@@
We mention that in quantum dot qubits the background disorder drifts in the sample changing the disorder landscape slowly over time, creating a charge noise. Typically, these drifts are slow and the noise is quasistatic, which is what our model for disorder assumes. This quasistatic approximation for the charge disorder is consistent with the fact that the associated noise manifests the well-known ’1/f frequency spectrum’ indicating that the noise s maximized at low frequencies (i.e. small values of f), where of course the quasistatic approximation is essentially exact. But such a drift, even if it is quasistatic, of course causes the practical problem that an already perfectly tuned experimental quantum dot system drifts out of its tuned parameter regime over time, and must be retuned. This is the key problem. Our machine learning technique developed in this work, when embedded into the system, if successful, would enable the retuning in an automated way without manual human intervention since all that is needed is a measurement of the charge stability diagram allowing the system to be retuned, which can be incorporated into the measurement and tuning electronics. The only necessary ingredient is the training of the algorithm using a sufficient amount of input stability diagrams obtained from measurements on the quantum dot circuits, as established in the current work using our simulations with different background disorder (which simulates the effective drift and noise).
@@
Further, though the use of higher resolution charge stability diagram input, our investigation seems to show that for smaller coupling constants the accuracy in predictions of the disorder can be increased as as much as necessary to fit the needs of experimentalists. We also show that a smaller $t_{ij}$ tends to increase the fidelity of the results in the other parameters. Additionally, our results show that the method continues to work as the number of quantum dots increases with little to no decrease in fidelity. While we do not envision using our method for a whole computational device (due to the scaling of computation in generating the training data), our technique likely could be scaled up to match the locality of a device to tune many quantum dots at once.
@@
Looking to the future, while our method is expected to be able to apply directly to experiment and improve the present tuning process of quantum dot devices, it should be possible to extend it to start from a microscopic Hamiltonian to determine directly the semiconductor quantum dot parameters. In fact, there is no reason to think that the technique itself is limited only to the Hubbard model, and in principle, it can be used to study other models of quantum dot qubits effectively. Our current use of the extended Hubbard model is based only on the fact that such a model seems to describe the semiconductor quantum dot based qubit platforms reasonably well.